Lady bug operates a small machine shop. He manufactures one standard product available from many other similar businesses, and he also manufactures products to customer order. His accountant prepared the annual income statement shown below:

Custom sales standard sales Total
Sales 50,000 25,000 75,000
Materials 10,000 8,000 18,000
Labor 20,000 9,000 29,000
Depreciation 6,300 3,600 9,900
power 700 400 1,100
rent 6,000 1,000 7,000
heat and light 600 100 700
other 400 900 1,300
total costs 44,000 23,000 67,000
Operating income 6,000 2,000 8,000

The depreciation charges are for machines used in the respective product lines. The power charge is apportioned on the estimate of power consumed. The rent is for the building space, which has been leased for 10 years at $7,000 per year. The rent and heat and light are apportioned to the product lines based on amount of floor space occupied. All other costs are current expenses identified with the product line incurring them.
A valued custom parts customer has asked Lady Bug to manufacture 5,000 special units for him. Lady Bug is working at capacity and would have to give up some other business to take this business. They cannot renege on custom orders already agreed to, but they could reduce the output of his standard product by about one-half for one year while producing the specially requested custom part. The customer is willing to pay $7 for each part. The materials cost will be about $2 per unit, and the labor will be $3.60 per unit. Lady Bug will have to spend $2,000 for a special device that will be discarded when the job is done.


The opportunity cost of taking the order is
A. $9,800 B. $3,550 C. $12,500 D. $4,000
What would be the cash advantage (disadvantage) if Lady Bug accepts the order?
A. $22,500 B. $13,950 C. $2,450 D. $1,450
Answer #1

1. Since depression, power , rent , heat & light expense are not affected by this project , they will not be considered for calculating opportunity cost.

Opportunity cost will be the contribution foregone because of this new offer.

Total relevent contribution = sales - material - labor - other expenses

= 25000 - 8000 - 9000 - 900

= 7100

Since half of the standard product will be foregone , hence opportunity cost = 7100 * 1/2 = $3550

Answer is option B

2. Cash advantage = contribution received from new offer - opportunity cost

Contribution Received = [($7 - $ 2 - $3.60)*5000] - $2000

= $5000

Cash advantage = $5000 - $3550

= $1450

Hence answer is option D.